OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, –4) with a slope of 1/2 y – 4 = 1/2 (x – 1) y + 4 = 2(x – 1) y + 4 = 1/2 (x – 1) y – 4 = 2(x + 1)

OpenStudy (hoblos):

the formula says: for a line passing through the point (a,b) and having slope m, its equation is of the form y-b=m(x-a) here a=1 , b=-4 & m=1/2 can you replace them ?

OpenStudy (anonymous):

Well, you use the formula y2-y1=M(x2-x1). You've got the points (1, -4), which we'll consider (x1, Y1), and a slope of 1/2, which is M. You simply plug them into the previous formula: y-(-4)=1/2(x-1)

OpenStudy (anonymous):

thank you guys. I think i've got it (: